OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 32x2 - 144

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

we don't even need to know that to find all the zeros

OpenStudy (misty1212):

first we solve \[u^2-32u-144=0\] then we replace \(u\) by \(x^2\)

OpenStudy (misty1212):

of course if you want to use that piece of information we can if \(-2i\) is a zero then so is \(2i\) so one of the factors is \[(x-2i)(x+2i)=x^2+4\]

OpenStudy (misty1212):

so it factors as \[(x^2+4)(x^2-36)\] and now the other zeros are easy to find

OpenStudy (anonymous):

Are they 0,-6,6?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

they are -6, 6 and 2i

OpenStudy (anonymous):

OK thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]